Sarge completed of his report in of an hour. At this rate, how much of the report will Sarge complete in 1 hour?
1
Expert's answer
2020-12-03T15:33:08-0500

Serge completed 1/10 of his report in 4/5 of an hour. At this rate, how much of the report will Serge complete in 1 hour?


"\frac{\frac{1}{10}}{x}=\frac{\frac45}{1}"

"x=\frac{1}{10}:\frac45=\frac18" of the report will be completed in 1 hour.